OpenStudy (anonymous):

d/dx * (integral of ln(t^2+1)dt from 0 to x^3) is? I will write it with the quation thing

OpenStudy (anonymous):

\[d/dx ( \int\limits_{0}^{x^3} \ln (t^2 +1)dt) \]

OpenStudy (anonymous):

don't even think about integrating that

OpenStudy (anonymous):

plug in x^3 for t

OpenStudy (anonymous):

\[d/dx \int\limits_{0}^{x^3} f(t )dt\] \[d/dx |_{0}^{x^3} F(t)dt\] \[d/dx ( F(x^3) - F(0) dt)=d/dx F(x^3)= 3x^2 f(x^3)\]

OpenStudy (anonymous):

\[f(t)=ln((t^2+1))\] f(x^3)=ln((x^6+1)) answer should be (from above post) \[3x^2 ln (x^6+1)\]

OpenStudy (anonymous):

thanks!!! =)

OpenStudy (anonymous):

it was really confusing me

OpenStudy (anonymous):

i was sittin trying to integrate XD

OpenStudy (anonymous):

Generally you see d/dx and integral , you don't have to integrate. Like for this problem, integral for kind of complicated if we integrate

OpenStudy (anonymous):

can i ask u a question about integrals?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when u have to integrate two functions that are being multiplied how shud u go about doing it? For example, \[\int\limits_{0}^{1} (x+1) \sqrt{x} dx \]

OpenStudy (anonymous):

for this particular problem you are better off distributing it \[(x+1)x^{1/2}=x^{3/2}+x^{1/2}\]